# Theory of Microphase Separation in Elastomers

### v1

@@ III Phase diagram
$\Gamma=b^{-1}\left[a(T-T_{\text{c}})+3b\psi_{0}^{2}+M\gamma^{-1}(1+\ln\gamma)\right].$ (S21)
@@
$f_{\text{U}}=\frac{1}{2}\left[a(T-T_{\text{c}})+M\right]\psi_{0}^{2}+\frac{1}{4}b\psi_{0}^{4}.$ (S22)
@@
For the stripe phase, we consider the stripe solution $\psi_{\text{S}}=\psi_{0}+A\cos(qx)$, with $A\cos(qx)$ representing a modulation of amplitude $A$ and wavenumber $q$ directed along one of the spatial directions. We substitute this solution in Eq. (S10) and minimize the free energy with respect to $q$ and $A$ to obtain the free-energy density $f_{\text{S}}$ of the stripe phase as
$f_{\text{S}}=f_{\text{U}}-\frac{b}{6}\Gamma^{2},\quad\text{with}\quad q_{\text{m}}^{2}=h^{-2}\ln\gamma\quad\text{and}\quad A^{2}=-\frac{4}{3}\Gamma.$ (S23)
As the modulation amplitude must be real for the stripe solution to exit, setting $\Gamma=0$ in Eq. (S21) gives an estimate for the microphase separation temperature $T_{\text{m}}$, which we find to be
$T_{\text{m}}=T_{\text{c}}-a^{-1}\left[3b\psi_{0}^{2}+M\gamma^{-1}\left(1+\ln\gamma\right)\right].$ (S24)
The same expression for $T_{\text{m}}$ can also be derived by linear stability analysis of the fluctuations around the uniform phase, as in Eq. (10) of the Letter.
@@
$q_{\text{m}}^{2}=h^{-2}\ln\gamma\quad\text{and}\quad A_{\pm}=\frac{4}{5}\left[\psi_{0}\pm\frac{1}{3}\left(9\psi_{0}^{2}-15\Gamma\right)^{1/2}\right].$ (S26)
The solution $A_{+}$ corresponds to the droplet phase, and $A_{-}$ corresponds to the “inverted” droplet phase, with the minimized free-energy density given by
$f_{\text{D}}=f_{\text{U}}-\frac{3b}{64}A_{\pm}^{2}\left(\psi_{0}A_{\pm}-2\Gamma\right).$ (S27)
@@
Above, the subscripts $i,j$ refer to one of the three phases: U, S, or D.
IV General coarse-graining procedure
@@
We now describe a general procedure to construct a large class of useful coarse-graining kernels in 3D. Although the procedure is not the only method, many routinely used kernels can be constructed using this method.
@@
where $\delta(x)$ is the Dirac delta function. We additionally assume that $P(x)$ has a well defined second moment $\mu_{2}=\int_{0}^{\infty}\mathrm{d}{x}\,x^{2}P(x)$ on the half-interval $[0,\infty)$. We can now use the density $P_{h}(x)$ to define a spherically-symmetric convolution kernel $K_{h}(\bm{x})$ in 3D as
@@
$F_{\bm{q}}=a(T-T_{\text{c}})+\kappa q^{2}+MK_{h}(\bm{q}).$ (S33)
The critical wavenumber $q_{\text{m}}$ is where $F_{\bm{q}}$ acquires its minimum. Using Eq. (S32), we find that $q_{\text{m}}$ satisfies the equation
$02\kappa q_{\text{m}}+MK_{h}^{\prime}(q_{\text{m}})=2\kappa q_{\text{m}}+MhK_{1}^{\prime}(hq_{\text{m}})=$ 0 . (S34)
@@
$02+\gamma\left[(hq_{\text{m}})^{-1}K_{1}^{\prime}(hq_{\text{m}})\right]=$ 0 . (S35)
The coarse-graining length scale $h\propto\xi\sim Y^{-1/2}$ using Eq. (S17) and the rescaled longitudinal modulus $M\sim Y$ from Eq. (S18). Because of this, we see that the parameter $\gamma$ is independent of both the Young’s modulus $Y$ and $h$, and Eq. (S35) would generally be a transcendental equation in the product $hq_{\text{m}}$. Any valid solution to this equation must always scale as $q_{\text{m}}\sim h^{-1}\sim Y^{1/2}$ and the product $hq_{\text{m}}$ would be a function of $\gamma$ alone. Therefore, irrespective of the kernel $K_{h}$, we see that the domain size $\Lambda$ follows the scaling
@@
The microphase separation temperature $T_{\text{m}}$ is the temperature at which the fluctuations around the uniform state become linearly unstable. As this occurs when $F_{\bm{q}}=-3b\psi_{0}^{2}$ and $q=q_{\text{m}}$, using Eqs. (S32) and (S33), we find
$T_{\text{m}}=T_{\text{c}}-a^{-1}\left[3b\psi_{0}^{2}+\kappa q_{\text{m}}^{2}+MK_{1}(hq_{\text{m}})\right].$ (S37)
Since the product $hq_{\text{m}}$ is a function of $\gamma$ alone, with $q_{\text{m}}^{2}\sim Y$ and $M\sim Y$, we see that $T_{\text{m}}$ always decreases linearly with $Y$, independent of the kernel $K_{h}$.
V Numerical techniques
As the field $\psi(\bm{x})$ is conserved (i.e., its spatial average is a constant), we can numerically minimize the free energy in Eq. (S10) by considering a “time”-dependent field [^6] We use Eq. (S38) solely for the purposes for energy minimization. It will not describe the actual dynamical evolution of $\psi$ seen in experiments, as it completely ignores the transport of material via hydrodynamic flow and related dissipative effects [28]. $\psi(\bm{x},t)$ and evolving it using Model B dynamics [61, 49]:
$\frac{\partial\psi(\bm{x},t)}{\partial t}=\nabla^{2}\left(\frac{\delta\mathscr{F}}{\delta\psi}\right)=\nabla^{2}\left[a(T-T_{\text{c}})\psi(\bm{x},t)+b\psi^{3}(\bm{x},t)-\kappa\nabla^{2}\psi(\bm{x},t)+M\bar{\psi}(\bm{x},t)\right].$ (S38)
Here $\nabla^{2}$ is the Laplacian, $t$ is the time, and $\bar{\psi}$ is the coarse-grained field defined in Eq. (S11). Time evolution is performed after Fourier transforming Eq. (S38) in space, which lets us compute spatial derivatives and convolutions efficiently using fast-Fourier techniques (assuming periodic boundary conditions). However, it is known that explicit first-order time stepping can lead to instabilities unless the time step $\delta t$ is very small. Therefore, we use a semi-implicit method employing a linearly stabilized splitting scheme [62, 63, 64] to evolve the time-discretized version of Eq. (S38) in Fourier space, given by

### v2

@@ III Phase diagram
$\Gamma=b^{-1}\left[a(T-T_{*})+3b\psi_{0}^{2}+M\gamma^{-1}(1+\ln\gamma)\right].$ (S24)
@@
$f_{\text{U}}=\frac{1}{2}\left[a(T-T_{*})+M\right]\psi_{0}^{2}+\frac{1}{4}b\psi_{0}^{4}.$ (S25)
@@
For the stripe phase, we consider the stripe solution $\psi_{\text{S}}=\psi_{0}+A\cos(qx)$, with $A\cos(qx)$ representing a modulation of amplitude $A$ and wavenumber $q$ directed along one of the spatial directions. We substitute this solution in Eq. (S10) and minimize the free energy with respect to $q$ and $A$ to obtain the free-energy density $f_{\text{S}}$ of the stripe phase at $q=q_{\text{m}}$ and $A=A_{\text{m}}$ as
$f_{\text{S}}=f_{\text{U}}-\frac{b}{6}\Gamma^{2},\quad\text{with}\quad q_{\text{m}}^{2}=h^{-2}\ln\gamma\quad\text{and}\quad A_{\text{m}}^{2}=-\frac{4}{3}\Gamma.$ (S26)
As the modulation amplitude must be real for the stripe solution to exit, setting $\Gamma=0$ in Eq. (S24) gives an estimate for the microphase separation temperature $T_{\text{micro}}$, which we find to be
$T_{\text{micro}}=T_{*}-a^{-1}\left[3b\psi_{0}^{2}+M\gamma^{-1}\left(1+\ln\gamma\right)\right].$ (S27)
The same expression for $T_{\text{micro}}$ can also be derived by linear stability analysis of the fluctuations around the uniform phase, as in Eq. (10) of the Letter.
@@
$q_{\text{m}}^{2}=h^{-2}\ln\gamma\quad\text{and}\quad A_{\text{m},\pm}=\frac{4}{5}\left[\psi_{0}\pm\frac{1}{3}\left(9\psi_{0}^{2}-15\Gamma\right)^{1/2}\right].$ (S29)
The solution $A_{\text{m},+}$ corresponds to the droplet phase, and $A_{\text{m},-}$ corresponds to the “inverted” droplet phase, with the minimized free-energy density given by
$f_{\text{D}}=f_{\text{U}}-\frac{3b}{64}A_{\text{m},\pm}^{2}\left(\psi_{0}A_{\text{m},\pm}-2\Gamma\right).$ (S30)
@@
Here, the subscripts $i,j$ refer to one of the three phases: U, S, or D.
A phase diagram constructed using the free-energy densities derived above for 2D modulations and choosing general, unitless parameters is presented in Fig. S2(a). Close to the critical point, three distinct phases can be observed: a uniform phase, a droplet (hexagonal) phase characterized by solvent-rich droplets dispersed within a solvent-deficient region, and a stripe phase consisting of alternating solvent-rich and solvent-deficient layers. Additionally, there is an inverted droplet phase consisting of solvent-deficient “droplets” embedded in a solvent-rich region. Compared to the phase diagram in Fig. 3 of the Letter, here we see the four phase coexistence regions more conspicuously. The general predictions of the phase diagram agree rather well with numerically obtained equilibrium morphologies in Figs. S2(b)–S2(e). In Fig. S2(a), all phase transitions are first order except at the critical point where a continuous transition from the uniform phase to the stripe phase is possible.
Intriguingly, the phase behavior of the system in 1D (Fig. S3) is somewhat distinct from the 2D situation. In 1D, the only modulated phase that one considers is a “patterned” phase $\psi(x)=\psi_{0}+A\cos(qx)$, equivalent to the stripe phase in 2D and 3D [53]. From Fig. S3, we see that in the temperature range immediately below the critical point, the system can transition from the uniform phase to the patterned phase continuously (indicated by a second-order transition curve. [^5] Similar second-order transition curves have also been predicted in the 1D version of the phase-field crystal model [96, 54] and in phenomenological models of membrane adhesion [55] analogous to the Blume–Emery–Griffiths model for 3He-4He mixtures [56].) Phase separation becomes first order below a tricritical point, leading to the emergence of a coexistence region between the uniform and patterned phases. The second-order transition curve in Fig. S3, given by $\Gamma=0$ [see Eq. (S24)], also exists in the 2D phase diagram of Fig. S2(a). However, it always falls within the phase boundaries of the other phases and for this reason, a continuous transition to the patterned (stripe) phase is not generally seen in 2D or 3D. Here it is worth noting that the recent study by Qiang et al. [31] on the strong-segregation limit of microphase separation in elastomers [31] also identifies a second-order phase transition at high stiffnesses and the presence of tricritical points. While their model differs from ours in several key aspects, the fact that their analysis is in 1D is the likely explanation for the observed continuous transition to patterned phases.
Figure S3: (a) General phase diagram in the $(\phi_{0},T)$ plane considering only 1D modulations for the same parameters as in Fig. S2. The dashed red curve indicates a second-order transition line, across which a continuous phase transition between the uniform (U) and the patterned (P) phases can occur. Solid blue curves are first-order transition lines. These curves meet at two tricritical points (marked with red dots). Regions of phase coexistence between the patterned and the uniform phases (P + U) are shaded in blue. (b),(c) Example equilibrium profiles $\psi(x)$ for various values of the mean order parameter $\psi_{0}$ (marked in the phase diagram).
Similarity to other pattern-forming systems
The phase diagram in Fig. S2 is rather similar to phase diagrams of other pattern-forming systems such as block copolymers and Langmuir monolayers [90]. Given this similarity it is worth examining the connection further. If we are primarily interested in modulations in $\psi$ with wavelengths much larger than the coarse-graining length scale, i.e., when $h^{2}q^{2}\ll 1$, we can expand $\bar{\psi}_{\bm{q}}=K_{h}(\bm{q})\psi_{\bm{q}}=\mathrm{e}^{-h^{2}q^{2}}\psi_{\bm{q}}$ in Eq. (S8) in powers of $q$ to find
$\bar{\psi}(\bm{q})=\left[1-h^{2}q^{2}+\frac{1}{2}h^{4}q^{4}+\cdots\right]\psi(\bm{q}).$ (S32)
Taking the inverse Fourier transform of the above expression, we identify the real-space representation of the coarse-grained order parameter $\bar{\psi}(\bm{x})$ to be
$\bar{\psi}(\bm{x})=\left[1+h^{2}\nabla^{2}+\frac{1}{2}h^{4}\nabla^{4}+\cdots\right]\psi(\bm{x}).$ (S33)
Substituting this in Eq. (S10) and after sufficient integration by parts to ensure that only powers of $\psi$ and $\nabla^{2}\psi$ remain, we find that the total free energy takes the form
$\mathscr{F}[\psi]\sim\int\mathrm{d}^{3}x\,\left\{\frac{1}{2}a(T-T_{**})\psi^{2}+\frac{1}{4}b\psi^{4}+\frac{1}{4}Mh^{4}\left[(\nabla^{2}+q_{\text{m}}^{2})\psi\right]^{2}-\eta\psi\right\},$ (S34)
where the wavenumber $q_{\text{m}}$ and the temperature $T_{**}$ are given by
$q_{\text{m}}^{2}=h^{-2}(1-\gamma^{-1})\quad\text{and}\quad T_{**}=T_{*}-a^{-1}M\left[1-\frac{1}{2}(1-\gamma^{-1})^{2}\right].$ (S35)
The above values agree to $\mathcal{O}[(\gamma-1)^{2}]$ with the corresponding expressions in Eq. (9) and Eq. (10) of the Letter in the limit $\gamma\to 1^{+}$. The free energy in Eq. (S34) is the Landau–Brazovskii free energy, which has been used to model a host of pattern-forming systems. Within the context of block copolymers, Fredrickson and Helfand [93] simplified Leibler’s free energy functional [92] for diblock copolymers to the above form. Apart from the phases in Fig. S2, a wide range of equilibrium morphologies, including BCC and gyroid phases, have been predicted for systems following this free energy in 3D [57].
IV Fluctuation effects, scattering, and Lifshitz behavior
To understand the effect of fluctuations above the transition temperature, we set $\psi=\psi_{0}+\delta\psi$ and expand Eq. (S8) to $\mathcal{O}(\delta\psi^{2})$, and find the quadratic (Gaussian) free energy of the fluctuations $\delta\psi$ in Fourier space as
$\mathscr{F}_{\text{G}}[\delta\psi]=\frac{1}{2}\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}(F_{\bm{q}}+3b\psi_{0}^{2})\,\delta\psi_{\bm{q}}\,\delta\psi_{-\bm{q}}.$ (S36)
The intensity distribution due to fluctuations seen in scattering experiments at equilibrium is proportional to the static structure factor $S(\bm{q})=\langle\delta\psi_{\bm{q}}\,\delta\psi_{-\bm{q}}\rangle$. For a quadratic free energy functional such as $\mathscr{F}_{\text{G}}$, it is given by [56]
$S(\bm{q})=\frac{\int\mathcal{D}[\delta\psi_{\bm{q}}]\,\mathcal{D}[\delta\psi_{-\bm{q}}]\,\delta\psi_{\bm{q}}\,\delta\psi_{-\bm{q}}\,\mathrm{e}^{-\mathscr{F}_{\text{G}}/(k_{\text{B}}T)}}{\int\mathcal{D}[\delta\psi_{\bm{q}}]\,\mathcal{D}[\delta\psi_{-\bm{q}}]\,\mathrm{e}^{-\mathscr{F}_{\text{G}}/(k_{\text{B}}T)}}$ $=\frac{k_{\text{B}}T}{F_{\bm{q}}+3b\psi_{0}^{2}}$ (S37)
$=\frac{k_{\text{B}}T}{a(T-T_{*})+3b\psi_{0}^{2}+\kappa q^{2}+M\mathrm{e}^{-h^{2}q^{2}}}.$ (S38)
For $M=0$, the structure factor has an Ornstein–Zernike form $S(q)\sim[1+(qr)^{2}]^{-1}$ with a mean-field correlation length $r\sim(T-T_{*})^{-1/2}$ as expected in general phase-separating systems. When $M=\phi_{*}^{-2}(\lambda+2\mu)\neq 0$ and $h=0$, i.e., when there is no mesoscopic length scale associated with elasticity, $S(\bm{q})$ would have the same form, but with a shifted critical temperature. The associated thermal correlation length close to the critical point is given by
$r^{2}=\frac{\kappa\phi_{*}^{2}}{a(T-T_{*})\phi_{*}^{2}+\lambda+2\mu}=\frac{\kappa\phi_{*}^{2}}{K_{\text{os}}+4\mu/3}.$ (S39)
In the last expression, we have written the denominator in terms of the osmotic bulk modulus $K_{\text{os}}=\phi_{*}^{2}f^{\prime\prime}(\phi_{*})+\lambda+2\mu/3$ noting that $f^{\prime\prime}(\phi_{*})=a(T-T_{*})$. In general, isotropic elastic systems become mechanically unstable when $K_{\text{os}}<0$. However, because the shear modulus $\mu$ is nonzero, we see that the correlation length $r$ in Eq. (S39) diverges only when $K_{\text{os}}\to-4\mu/3$. This is a well known result, particularly in the context of gels [28] where a distinction is made between the mechanical instability point given by $K_{\text{os}}=0$, and the “cloud point” given by $K_{\text{os}}+4\mu/3=0$ where the gel loses thermodynamic stability. However, macroscopic shape changes in gels occur over a significant amount of time and the mechanical instability at $K_{\text{os}}=0$ can be hard to observe unless the gel sample is very small in size [58, 70]. For this reason, gels can remain stable even when $K_{\text{os}}<0$ and can exhibit anomalous properties like a negative Poisson’s ratio when $K_{\text{os}}<2\mu/3$ [59]. At temperatures below the cloud point, the gel becomes opaque and undergoes spinodal decomposition [25, 50].
[…]
When elasticity operates at a nonzero mesoscopic length scale $h$ and the parameter $\gamma=Mh^{2}/\kappa>1$, the structure factor has a peak at $q_{\text{m}}^{2}=h^{2}\ln\gamma$. Furthermore, the peak value $S(q_{\text{m}})\sim(T-T_{\text{micro}})^{-1}$, with $T_{\text{micro}}$ being the microphase separation temperature defined in Eq. (10) of the Letter. An example $S(\bm{q})$ for the experimental parameter ranges is illustrated in Fig. S4. When the elastomer is cooled from $T\to T_{\text{micro}}$, we expect the scattering intensity to have a smoothly growing peak at a fixed $q$ similar to $S(\bm{q})$, in agreement with the experiments of Ref. [24]. In the context of crosslinked polymer blends, de Gennes’s prediction of $S(0)=0$ [100] was not observed experimentally [98], and lead to other theoretical efforts [99] to explain the discrepancy. Interestingly, our model predicts $S(0)\neq 0$. This, along with the similarity in the scaling behaviors of the domain size and the transition temperature, strongly suggests that crosslinked polymer blends may also have an internal elastic response that is nonlocal.
To glean more analytical insights into the general nature of the long-range correlations associated with the structure factor in Eq. (S38), assuming small $q$, we expand the denominator to $\mathcal{O}(q^{4})$ to find
$S(\bm{q})\sim\frac{S(0)\,\tau}{\tau+2(1-\gamma)q^{2}+\gamma h^{2}q^{4}},\quad\text{where}\quad\tau=2\kappa^{-1}[a(T-T_{*})+3b\psi_{0}^{2}+M].$ (S40)
This structure factor can also be derived from the simplified free energy in Eq. (S34) and resembles the structure factor observed in microemulsions—fluid mixtures of water and oil stabilized by a surfactant [104]. The Fourier inverse of the structure factor $S(\bm{q})$ is the real-space two-point correlation function $C(\bm{x})=\langle\delta\psi(\bm{x})\delta\psi(0)\rangle$. Intuitively, $C(\bm{x})$ is a measure of the probability of the values of $\delta\psi$ at a point $\bm{x}$ and at the origin being the same.
For clarity in presenting the results below, we define the following nondimensional parameter:
$\zeta=\frac{(1-\gamma)}{\sqrt{\gamma\tau h^{2}}}.$ (S41)
When $\left|\zeta\right|<1$, the correlation function associated with Eq. (S40) is of the form [104, 105]
$C(\bm{x})=C(0)\,\mathrm{e}^{-\left|\bm{x}\right|/r}\operatorname{sinc}\left(\frac{2\pi\left|\bm{x}\right|}{d}\right),$ (S42)
and represents exponentially decaying oscillations with a period $d$ and correlation length $r$, given by
$r=\left(\frac{\gamma h^{2}}{\tau}\right)^{1/4}\sqrt{\frac{2}{1+\zeta}}\quad\text{and}\quad d=\left(\frac{\gamma h^{2}}{\tau}\right)^{1/4}\sqrt{\frac{2}{1-\zeta}}.$ (S43)
In microemulsions, a parameter analogous to $\zeta$ measures the strength of the surfactant: a stronger surfactant results in a negative $\zeta$ and vice versa [105]. For elastomers, we see from Eq. (S41) that $\zeta$ is controlled by the parameter $\gamma=Mh^{2}/\kappa$, which is analogous to the (inverse) elastocapillary number, and measures the relative importance of the elastic and surface contributions to the free energy: $\gamma>1$ results in $\zeta<0$ and vice versa. The phase behavior of elastomers above their transition temperature can also be compared to that of microemulsions, as outlined below and graphically represented in Fig. S5.
  • When $\left|\zeta\right|<1$, both $r$ and $d$ are finite and we consider the system to be in a structured-disordered phase (also called the “middle” or the bicontinuous microemulsion BµE phase), referring to the presence of fluctuating mesoscopic structures within an overall disordered system [105]. This is indicated by a correlation function with decaying oscillations.
  • When $0<\zeta<1$ (which requires $\gamma<1$), the structure factor has a peak only at $q_{\text{m}}=0$, even though $C(\bm{x})$ continues to have an oscillatory behavior. As $\zeta\to 1$, the oscillation period $d$ diverges and the curve where $\zeta=1$ is the disorder line [105].
    For $\zeta>1$, the correlation function $C(\bm{x})$ decays monotonically without oscillations (disordered phase). When $\zeta\gg 1$, the coefficient of $q^{2}$ in the denominator of Eq. (S40) is large and positive and the $q^{4}$ term can be neglected. This leads to an Ornstein–Zernike-like $S(\bm{q})$, with a correlation function that decays exponentially.
  • The line where $\zeta=0$ (or $\gamma=1$) is the Lifshitz line [77]. Beyond the Lifshitz line, when $\gamma>1$ and $1<\zeta<0$, the structure factor has a peak at $q_{\text{m}}^{2}=h^{-2}(1-\gamma^{-1})$ signifying the dominant presence of fluctuating structures. When $\zeta\to-1$, the correlation length $r$ diverges, indicating spatial ordering and the emergence of an ordered phase. Microphases appear for $\zeta<-1$.
Finally, it should be emphasized that in the presence of fluctuations, the mean-field phase diagrams we have considered accurately describe the phase behavior only when sufficiently far from the critical point. A more complete analysis of the effect of fluctuations for systems described by the free-energy functional of Eq. (S34) was presented by Brazovskii using a Hartree approximation [101]. Fluctuations were observed to suppress the second-order phase transition at the critical point, which gets replaced by a first-order transition line across which the system can transition from a uniform phase to a stripe phase. Away from the critical point, fluctuations effectively lead to specific compositional ranges of the mean order parameter $\psi_{0}$ (or $\phi_{0}$) that allow for a direct transition from the uniform phase to various ordered phases as temperature changes. For block copolymers, fluctuation effects can be significant [60, 61]. Since the free energy in our model can be simplified to the form in Eq. (S34), we anticipate that fluctuations will influence the phase behavior very close to the order-disorder transition temperature. For example, with the inclusion of fluctuations, a direct transition between the disordered and gyroid phases may occur [62], which could, for example, explain the experimental observation of the gyroid-like bicontinuous phase for elastomers at high stiffnesses [24].
Figure S5: General phase diagram in the $(\tau,\gamma)$ plane with $h=1$. Here $\tau$ is an effective temperature and $\gamma$ is a dimensionless parameter analogous to the (inverse) elastocapillary number. See Eq. (S40) for definitions. The dashed red curve given by $\zeta=-1$ ($\gamma>1$) and the dashed red line given by $\tau=0$ ($\gamma<1$) indicate second-order transitions. A first-order triple line (solid blue curve) given by $\tau=-(2+\sqrt{6})(1-\gamma)^{2}/(h^{2}\gamma)$ separates the two-phase coexistence region from the stripe phase. The gray curve separating the bicontinuous microemulsion (BµE) phase with $q_{\text{m}}=0$ and the disordered one is the disorder line where $\zeta=1$. The gray vertical line where $\zeta=0$ $(\gamma=1)$ represents the Lifshitz line which ends at the Lifshitz point (black dot) [77]. Neither the disorder line nor the Lifshitz line indicate phase transitions.
V General coarse-graining procedure
@@
We now describe a general procedure to construct a large class of useful coarse-graining kernels in 3D. Although this procedure is not the only method, many routinely used kernels can be constructed using this method.
@@
where $\delta(x)$ is the Dirac delta function. We additionally assume that $P(x)$ has a well defined second moment $\mu_{2}=\int_{0}^{\infty}\mathrm{d}{x}\,x^{2}P(x)$ on the half-interval $0,\infty)$. We can now use the density $P_{h}(x)$ to define a spherically-symmetric convolution kernel $K_{h}(\bm{x})$ in 3D as
@@
$F_{\bm{q}}=a(T-T_{*})+\kappa q^{2}+MK_{h}(\bm{q}).$ (S48)
The wavenumber $q_{\text{m}}$ is where $F_{\bm{q}}$ acquires its minimum. Using Eq. (S47), we find that $q_{\text{m}}$ satisfies the equation
$2\kappa q_{\text{m}}+MK_{h}^{\prime}(q_{\text{m}})=2\kappa q_{\text{m}}+MhK_{1}^{\prime}(hq_{\text{m}})=0.$ (S49)
@@
$2+\gamma\left[(hq_{\text{m}})^{-1}K_{1}^{\prime}(hq_{\text{m}})\right]=0.$ (S50)
The coarse-graining length scale $h\sim\xi\sim Y^{-1/2}$ using Eq. (S21) and the rescaled longitudinal modulus $M\sim Y$ from Eq. (S18). Because of this, we see that the parameter $\gamma$ is independent of both the Young’s modulus $Y$ and $h$, and Eq. (S50) would generally be a transcendental equation in the product $hq_{\text{m}}$. Any valid solution to this equation must always scale as $q_{\text{m}}\sim h^{-1}\sim Y^{1/2}$ and the product $hq_{\text{m}}$ would be a function of $\gamma$ alone. Therefore, irrespective of the kernel $K_{h}$, we see that the domain size $\Lambda$ follows the scaling
@@
The microphase separation temperature $T_{\text{micro}}$ is the temperature at which the fluctuations around the uniform state become linearly unstable. As this occurs when $F_{\bm{q}}=-3b\psi_{0}^{2}$ and $q=q_{\text{m}}$, using Eqs. (S47) and (S48), we find
$T_{\text{micro}}=T_{*}-a^{-1}\left[3b\psi_{0}^{2}+\kappa q_{\text{m}}^{2}+MK_{1}(hq_{\text{m}})\right].$ (S52)
Since the product $hq_{\text{m}}$ is a function of $\gamma$ alone, with $q_{\text{m}}^{2}\sim Y$ and $M\sim Y$, we see that $T_{\text{micro}}$ always decreases linearly with $Y$, independent of the kernel $K_{h}$. The above two equations generalize the scaling behavior of $T_{\text{micro}}$ and $\Lambda$ presented in Eqs. (10) and (12) of the Letter for a Gaussian kernel.
VI Numerical techniques
As the field $\psi(\bm{x})$ is conserved (i.e., its spatial average is a constant), we can numerically minimize the free energy in Eq. (S10) by considering a “time”-dependent field $\psi(\bm{x},t)$ and evolving it using Model B dynamics [64, 97]:
$\frac{\partial\psi(\bm{x},t)}{\partial t}=\nabla^{2}\left(\frac{\delta\mathscr{F}}{\delta\psi}\right)=\nabla^{2}\left[a(T-T_{*})\psi(\bm{x},t)+b\psi^{3}(\bm{x},t)-\kappa\nabla^{2}\psi(\bm{x},t)+M\bar{\psi}(\bm{x},t)\right].$ (S53)
Here $\nabla^{2}$ is the Laplacian, $t$ is the time, and $\bar{\psi}$ is the coarse-grained field defined in Eq. (S11). We use Eq. (S53) solely for the purposes of energy minimization. It will not describe the actual dynamical evolution of $\psi$ seen in experiments, as it completely ignores material transport via hydrodynamic flow and related dissipative effects [41].
Time evolution is performed after Fourier transforming Eq. (S53) in space, which lets us compute spatial derivatives and convolutions efficiently using fast-Fourier techniques (assuming periodic boundary conditions). However, it is known that explicit first-order time stepping can lead to instabilities unless the time step $\delta t$ is very small. Therefore, we use a semi-implicit method employing a linearly stabilized splitting scheme [65, 66, 67] to evolve the time-discretized version of Eq. (S53) in Fourier space, given by